OpenStudy (tiffanymak1996):

Find limit:

OpenStudy (bahrom7893):

of?

OpenStudy (tiffanymak1996):

\[\lim_{(x,y) \rightarrow (0,0)} \frac{ x^2y^3}{ \sqrt(x^2 +y^3) } \]

OpenStudy (tiffanymak1996):

evaluate the limit by converting to polar coordinates

OpenStudy (bahrom7893):

to convert into polar coordinates, use: \[x=r*Cos(\theta)\]\[y=r*Sin(\theta)\]\[x^2+y^2=(r*Cos(\theta))^2+(r*Sin(\theta))^2=r^2Cos^2(\theta)+r^2Sin^2(\theta)\]\[x^2+y^2=r^2(Cos^2(\theta)+Sin^2(\theta))=r^2*1=r^2\]

OpenStudy (tiffanymak1996):

I know how to convert it, but I'm a little stuck with the theta involved in the limit.

OpenStudy (bahrom7893):

So that whole limit becomes:\[(r=\sqrt{x^2+y^2}=\sqrt{0^2+0^2}=0)\] \[\lim_{r \rightarrow 0} \frac{ r^2Cos^2(\theta)r^3Sin^3(\theta)}{ \sqrt{r^2Cos^2(\theta) +r^3Sin^3(\theta)} }\]

OpenStudy (bahrom7893):

Okay, I'm stuck. Was reading this, if it's of any help: http://www.maths.adelaide.edu.au/michael.albanese/other/polar.pdf @amistre64 mind helping us out here?

OpenStudy (tiffanymak1996):

I got stuck when i reach this: \[\lim_{r \rightarrow 0+} \frac{ r^5 \sin^22\theta \sin \theta }{ \sqrt(r(r+\sin \theta)) }\] I tried expanding the sin sq and sin using product to sum, but it didn't help much.

OpenStudy (amistre64):

\[\lim_{r \to 0} \frac{ r^2\cos^2(t)r^3\sin^3(t)}{ \sqrt{r^2\cos^2(t) +r^3\sin^3(t)} }\] \[\lim_{r \to 0} \frac{ r^2\cos^2(t)r^3\sin^3(t)}{r \sqrt{\cos^2(t) +r\sin^3(t)} }\] \[\lim_{r \to 0} \frac{ r^4\cos^2(t)\sin^3(t)}{\sqrt{\cos^2(t) +r\sin^3(t)} }\] seems to me that just zeros out the top without hurting the bottom

OpenStudy (tiffanymak1996):

ok, thanks!

OpenStudy (amistre64):

i cant say how correct it is, but that does seem to play out like that

OpenStudy (tiffanymak1996):

I have the answer in the book, it's 0, i just needed to know why and how it is zero.

OpenStudy (amistre64):

the offending part was the r attached to the underside .... by factoring out an r^2 we were able to remove its offensive behaviour

OpenStudy (tiffanymak1996):

@bahrom7893 thanks for helping!

OpenStudy (tiffanymak1996):

make sense, thanks!